# Multi-centre normative brain mapping of intracranial EEG lifespan patterns in the human brain

### v1

@@ Abstract
Background: Understanding healthy human brain function is crucial to identify and map pathological tissue within it. Whilst previous studies have mapped intracranial EEG (icEEG) from non-epileptogenic brain regions, these maps do not consider the effects of age and sex. Further, most existing work on icEEG has often suffered from a small sample size due to this modality’s invasive nature. Here, we substantially increase the subject sample size compared to existing literature, to create a multi-centre, normative map of brain activity which additionally considers the effects of age, sex and recording site.
Methods: Using interictal icEEG recordings from $n=513$ subjects originating from 15 centres, we constructed a normative map of non-pathological brain activity by regressing age and sex on relative band power in five frequency bands, whilst accounting for the site effect.
Results: Recording site significantly impacted normative icEEG maps in all frequency bands, and age was a more influential predictor of band power than sex. The age effect varied by frequency band, but no spatial patterns were observed at the region-specific level. Certainty about regression coefficients was also frequency band specific and moderately impacted by sample size.
Conclusion: The concept of a normative map is well-established in neuroscience research and particularly relevant to the icEEG modality, which does not allow healthy control baselines. Our key results regarding the site and age effect guide future work utilising normative maps in icEEG.
@@ 1 Introduction
To analyse this effect thoroughly, this investigation utilises a large data set, encompassing multiple international hospital sites. This is similar to Bethlehem et al. (2022) which leverages MRI scans from various studies globally to create brain charts for the human lifespan. Importantly, we will perform our analysis on the largest multi-centre normative icEEG dataset ($n=513$ subjects) to date, accounting for site effects in data by using mixed-effect modelling. Following this, we aim to uncover and discuss the relationships, if any, between band power extracted from icEEG recordings, and the variables age, sex and site.
Ultimately, we highlight the need to account for the heterogeneity of icEEG data. We seek to fill a gap in the literature regarding the effect of two standard covariates – age and sex – on band power values extracted from icEEG and modelled in a normative setting. We also expand on previous sample sizes and properly consider site variability. Such research is necessary, as an understanding of ageing patterns, sex differences and site variation in this modality will conceivably aid in the identification of pathological activity through deviations from ‘normal’ trends.
@@ 2.1 Subjects
Our analysis involved 513 individuals with epilepsy undergoing presurgical evaluation with icEEG to localise the seizure onset zone. Data was collected from Beijing Tiantan Hospital, Great Ormond Street Hospital, University of Iowa Hospital, SickKids, University College London Hospital and the University of Wisconsin-Madison. We also used 9 sites from the publicly available RAM data set (https://memory.psych.upenn.edu/RAM), bringing our total number of data sites to 15.
Both grid and depth electrode recordings were included. A summary of our final cohort is provided in Figure 1.
Figure 1: The number of subjects from each participating site, along with their age and sex distributions. Sites from the RAM cohort are highlighted with an *.
@@ 2.2 Electrode localisation
The methods for localising the electrode contacts to brain regions have also been described previously (Taylor et al., 2022, Wang et al., 2020, Hamilton et al., 2017, Wang et al., 2023). Different sites provided different levels of data, so our methods for localising electrode contacts to ROIs varied between them. Sites either provided contact locations in MNI space, or provided native space imaging and co-localised contacts. In the first case, we assigned electrodes to one of 82 regions from the Lausanne scale 36 atlas (Hagmann et al., 2008). We used FreeSurfer to generate volumetric parcellations of an MNI space template brain (Fischl, 2012, Hagmann et al., 2008). Each electrode contact was assigned to the closest grey matter volumetric region within 5 mm. If the closest grey matter region was $>$5 mm away then the contact was excluded from further analysis. For the latter case, a similar technique was used, but applied in native space using the subject’s own parcellated pre-operative MRI.
To ensure our findings were robust to parcellation choice, we confirmed that key results held using a finer-grained parcellation (see Supplementary S3).
@@
2.3.1 Segment selection
For each subject, we extracted a 70-second interictal icEEG segment from a period of relaxed wakefulness, at least two hours away from any detected ictal events. The raw signals and the power spectral densities were visually and algorithmically inspected for artefacts and faulty channels. Using clinical reports and/or post-operative data, we excluded any contacts that: had frequent spikes, were within known lesions, were within the seizure onset zone, or were subsequently resected (if such information was available). Ensuring that at the subject level, we only retained channels thought to be non-pathological in terms of both location in the brain and the signal produced.
2.3.2 Signal processing
@@
In the $\gamma$-band, data between 47.5-52.5 Hz and 57.5-62.5 Hz were excluded from all sites, to avoid any power line noise. The $\gamma$-band was also capped at 77.5 Hz due to 80 Hz noise in the RAM sites. Band power estimates were log10 transformed and normalised to sum to one in each contact (L1-normalised). These final values are used throughout results to represent relative band power, denoted RBP($\cdot$).
@@ 2.4 Normative data table creation
Previous work has demonstrated a left/right symmetry in relative band power in EEG, MEG, and icEEG (Taylor et al., 2022, Janiukstyte et al., 2023, Owen et al., 2023b). Hence, to maximise the number of subjects in each region, we created a ‘mirrored’ version of our data table, in which, the data from homologous regions were combined (e.g., the left amygdala and the right amygdala). In the case of bilateral implantation of symmetric regions, we ensured individuals only had one value of RBP($\cdot$) per ROI and frequency band. This mirrored data table only comprised 38 regions but had markedly more subjects in each. Throughout the results, it will be made clear whether the mirrored, or original, normative data table is in use at any given time. We confirm the validity of this mirroring in Supplementary S2 by repeating one of our key results using the original data.
In summary, our final normative data table included a unique subject identifier, their age and sex, their originating site, and, for the regions in which they had electrode implantation only, their RBP($\cdot$) in five frequency bands averaged across contacts where necessary.
@@ 2.5 Statistical modelling and testing
The final step was to fit a suitable model to the normative data to examine the effect of age, sex and site on relative band power. Visual inspection of the scatter plots of the RBP($\cdot$) values against age in each frequency band (at the whole-brain level) showed no evidence of a non-linear relationship (see Supplementary S1). Further, the effect of having multiple recording sites had to be considered, especially as some sites supplied only paediatric, or only adult recordings. Therefore, we implemented a linear mixed model (LMM) in each frequency band, specifically a random intercept model, with the originating site as a random effect. Possible fixed effects were age and sex, meaning four LMMs were under consideration in each frequency band:
Null: $RBP(\cdot)\sim(1|Site)$
Age: $RBP(\cdot)\sim Age+(1|Site)$
Sex: $RBP(\cdot)\sim Sex+(1|Site)$
Full: $RBP(\cdot)\sim Age+Sex+(1|Site)$
Model fitting could be performed at either the regional or whole-brain level. Taking the latter, we calculated several model evaluation statistics to determine the optimal fixed effect structure of the LMM in each frequency band. Since we worked at the whole-brain level, the original data table was used. The statistics under consideration were the AIC and BIC for each model, 95% profiled confidence intervals for the regression coefficients of fixed effects in all but the null model (which has none) and the likelihood ratio test p-values for every pair of models which differed by one variable.
In every frequency band, each metric was in agreement regarding the optimal LMM, so no further tests were required. The preferred models were the age model in the $\delta$, $\alpha$ and $\beta$ bands; the sex model in $\gamma$ and the full model in $\theta$. Hence, the optimal LMM was frequency band-specific in our cohort, and so we report frequency band-specific effects throughout our results.
We have provided a brief visual overview of our methods (Figure 2). In the results section, we explore both random (site) and fixed (age, sex) effects. Specifically, we investigate if the variables explain RBP($\cdot$) variation, and examine how they might influence it.
Figure 2: A visual representation of our methods: electrode localisation (top) and signal processing (middle) following segment selection, for each subject. Subjects were combined to create our normative data set, which we applied our statistical models to (bottom). All components show dummy data for example purposes only. Brain plots from Scholtens et al. (2021).
@@ 3 Results
3.1 Site effects impacted icEEG normative maps
Focusing first on the random effect structure (recording hospital site), we found that in our cohort and model, the site effect was much more powerful in explaining RBP($\cdot$) differences at the whole brain level than any fixed effect structure, in almost all cases. Table 1 quantifies this, showing marginal $R^{2}$ values ($R^{2}_{m}$) and intraclass correlation coefficients ($ICC$) as percentages for the full model, the age model and the sex model. The $R^{2}_{m}$ represents the proportion of variation in RBP($\cdot$) explained by the fixed effect(s) of that model. The $ICC$ represents the proportion of the variance explained by the recording sites. The optimal LMM as per Section 2.5 has been highlighted for each frequency band. Note, the $R_{m}^{2}$ values of the two single fixed effect models do not sum exactly to the $R^{2}_{m}$ of the full model, because the denominator in the calculation of $R_{m}^{2}$ includes the residual variance and therefore changes with each model.
In all bands except $\alpha$, the $ICC$ was consistently larger than the $R^{2}_{m}$ values across models, with this difference being most notable in the $\theta$ and $\gamma$ bands. Whilst there is a range of $ICC$s in Table 1, they always exceed 5%, indicating that, in our cohort, at least a twentieth of the variation in RBP($\cdot$) was explained by site differences. This proportion sometimes reached as high as one-quarter, attributing that site impacted RBP($\cdot$) in all frequency bands.
@@ 3.1 Site effects impacted icEEG normative maps
$\delta$ 4.82 9.24 4.82 9.24 0.03 13.00
$\theta$ 1.27 11.26 0.79 11.15 0.62 13.07
$\alpha$ 7.40 5.87 7.38 5.88 0.02 6.22
$\beta$ 3.18 5.31 3.18 5.31 0.02 6.11
$\gamma$ 0.36 24.55 0.06 24.31 0.32 25.02
Table 1: $R^{2}_{m}$ and $ICC$ values (measured in %) for the full model, the age model and the sex model. $R^{2}_{m}$ represents the proportion of variation in RBP($\cdot$) explained by the fixed effect(s) of that model. The $ICC$ represents the proportion of the variance explained by the grouping structure, namely recording sites. The optimal covariate subset, as determined by a standard model selection process, is highlighted for each frequency band.
@@ 3.2 Age and sex explained only some of the variation in relative band power
We next examined our fixed effects at the whole-brain level ($R^{2}_{m}$). Neither age nor sex explained a notable portion of the variance in the RBP($\cdot$), in any frequency band (Table 1). The largest effect was found in the $\alpha$-band, where $R^{2}_{m}$ was highest at 7%, which in contrast, is similar to the minimum $ICC$ values at around 5%.
@@ 3.3 Age was more important than sex for predicting relative band power
Comparing $R^{2}_{m}$ for the age and sex models revealed which of the two was more valuable when predicting RBP($\cdot$). Age outperformed sex in all frequency bands except $\gamma$ (Table 1). In fact, sex consistently accounted for a negligible portion of the variance in the response, with $R^{2}_{m}$ values below 1%. This statement holds even in $\theta$ and $\gamma$, with sex retained in the optimal LMM, confirming sex was not a significant predictor in this cohort.
In contrast, $R^{2}_{m}$ values for age fluctuated notably across frequency bands, being higher in $\delta$, $\alpha$ and $\beta$ ($R^{2}_{m}$ range of $7.33\%$ and $0.61\%$ for age and sex models respectively). The frequency bands which retain the age model as the optimal LMM ($\delta$, $\alpha$, $\beta$) have the highest proportion of variance explained by the optimal fixed effects structure, emphasizing that age was the more important predictor of RBP($\cdot$) and the influence of sex was minor.
Subsequent analysis focuses on the age model for all frequency bands; however, we exercise caution when interpreting results in the $\theta$ and $\gamma$ bands.

### v2

@@ Abstract
Background: Understanding healthy human brain function is crucial to identify and map pathological tissue within it. Whilst previous studies have mapped intracranial EEG (icEEG) from non-epileptogenic brain regions, these maps do not consider the effects of age and sex. Further, most existing work on icEEG has often suffered from a small sample size due to the modality’s invasive nature. Here, we substantially increase the subject sample size compared to existing literature, to create a multi-centre, normative map of brain activity which additionally considers the effects of age, sex and recording hospital.
Methods: Using interictal icEEG recordings from $n=502$ subjects originating from 15 centres, we constructed a normative map of non-pathological brain activity by regressing age and sex on relative band power in five frequency bands, whilst accounting for the hospital effect.
Results: Recording hospital significantly impacted normative icEEG maps in all frequency bands, and age was a more influential predictor of band power than sex. The age effect varied by frequency band, but no spatial patterns were observed at the region-specific level. Certainty about regression coefficients was also frequency band specific and moderately impacted by sample size.
Conclusion: The concept of a normative map is well-established in neuroscience research and particularly relevant to the icEEG modality, which does not allow healthy control baselines. Our key results regarding the hospital site and age effect guide future work utilising normative maps in icEEG.
KEYWORDS: intracranial EEG, normative modelling, lifespan patterns, sex differences, hospital effects, band power
@@ 1 Introduction
To analyse this effect thoroughly, this investigation utilises a large data set, encompassing multiple international hospitals. This is similar to Bethlehem et al. (2022) which leverages MRI scans from various studies globally to create brain charts for the human lifespan – our study seeks to help establish such methods in the icEEG literature. Importantly, we will perform our analysis on the largest multi-centre normative icEEG dataset to date ($n=502$ subjects), accounting for hospital site effects in data by using mixed-effect modelling. Following this, we aim to uncover and discuss the relationships, if any, between band power extracted from icEEG recordings, and the variables age, sex and recording hospital.
Ultimately, we highlight the need to account for the heterogeneity of icEEG data. We seek to fill a gap in the literature regarding the effect of two standard covariates – age and sex – on band power values extracted from icEEG and modelled in a normative setting. We also expand on previous sample sizes and properly consider hospital variability. Such research is necessary, as an understanding of ageing patterns, sex differences and hospital variation in this modality will conceivably aid in the identification of pathological activity through deviations from ‘normal’ trends.
@@ 2.1 Subjects
Our analysis involved 502 individuals with epilepsy undergoing presurgical evaluation with icEEG to localise the seizure onset zone. Data was collected from Beijing Tiantan Hospital, Great Ormond Street Hospital, University of Iowa Hospital, SickKids, University College London Hospital and the University of Wisconsin-Madison. We also used nine hospitals from the publicly available RAM data set (https://memory.psych.upenn.edu/RAM), bringing our total number of contributing hospital sites to 15.
Anonymised icEEG recordings were analysed following approval of the Newcastle University Ethics Committee (reference number 23973/2022). Both grid and depth electrode recordings were included. A summary of our final cohort is provided in Figure 1.
Figure 1: The number of subjects from each participating hospital, along with age and sex distributions. Hospitals from the RAM cohort are highlighted with an *.
@@ 2.2 Electrode localisation
The methods for localising the electrode contacts to brain regions have also been described previously (Taylor et al., 2022, Wang et al., 2020, Hamilton et al., 2017, Wang et al., 2023). Different hospital sites provided different levels of data, so our methods for localising electrode contacts to ROIs varied between them. Hospitals either provided contact locations in MNI space, or provided native space imaging and co-localised contacts. In the first case, we assigned electrodes to one of 82 regions from the Lausanne scale 36 atlas (Hagmann et al., 2008). We used FreeSurfer to generate volumetric parcellations of an MNI space template brain (Fischl, 2012, Hagmann et al., 2008). Each electrode contact was assigned to the closest grey matter volumetric region within 5 mm. If the closest grey matter region was $>$5 mm away then the contact was excluded from further analysis. For the latter case, a similar technique was used, but applied in native space using the subject’s own parcellated pre-operative MRI.
To ensure our findings were robust to parcellation choice, we confirmed that key results held using a finer-grained parcellation (Supplementary S3).
@@
Segment selection
For each subject, we extracted a 70-second interictal icEEG segment from a period of relaxed wakefulness, at least two hours away from any detected ictal events. The raw signals and the power spectral densities were visually and algorithmically inspected for spikes, artefacts and faulty channels (see Supplementary S1 for details). Using clinical reports (if available) we excluded any contacts that were: within known lesions, within the seizure onset zone, or subsequently resected. This ensured that at the subject level, we only retained channels thought to be non-pathological in terms of both location in the brain and the signal produced.
If unavailable in clinical reports, information on resected contacts could also be obtained by drawing masks using post- and pre-operative scans (where available), as in Taylor et al. (2022) and Wang et al. (2020). Supplementary S2 details the exclusionary information available for each hospital.
Signal processing
@@
In the $\gamma$-band, data between 47.5-52.5 Hz and 57.5-62.5 Hz were excluded from all hospital sites, to avoid any power line noise. The $\gamma$-band was also capped at 77.5 Hz due to 80 Hz noise in the RAM cohort. Band power estimates were log10 transformed and normalised to sum to one in each contact (L1-normalised). These final values are used throughout results to represent relative band power, denoted RBP($\cdot$).
@@ 2.4 Normative data table creation
Previous work has demonstrated a left/right symmetry in RBP($\cdot$) in EEG, MEG, and icEEG (Taylor et al., 2022, Janiukstyte et al., 2023, Owen et al., 2023b). Hence, to maximise the number of subjects in each region, we created a ‘mirrored’ version of our data table, in which, the data from homologous regions were combined (e.g., the left amygdala and the right amygdala). In the case of bilateral implantation of symmetric regions, we ensured individuals only had one value of RBP($\cdot$) per ROI and frequency band. This mirrored data table only comprised 38 regions but had markedly more subjects in each. Throughout the results, it will be made clear whether the mirrored (regional), or original (whole-brain), normative data table is in use at any given time.
We confirm the validity of this mirroring in Supplementary S4 by repeating one of our key results using the original data. Additionally, Supplementary S5 demonstrates that our normative values are not related to a subject’s age of epilepsy onset, a feature representing pathology. Unfortunately, we do not have the data to check other disorder-specific features, such as drug levels or epilepsy classification.
In summary, our final normative data table included a unique subject identifier, their age and sex, their originating hospital, and, for the regions in which they had electrode implantation only, their RBP($\cdot$) in five frequency bands averaged across contacts where necessary.
@@ 2.5 Statistical modelling and testing
The final step was to fit a suitable model to the normative data to examine the effect of age, sex and hospital site on relative band power. Visual inspection of scatter plots of the RBP($\cdot$) values against age in each frequency band (at the whole-brain level) showed no evidence of a non-linear relationship (see Supplementary S6). Further, the effect of having multiple recording hospitals had to be considered, especially as some hospitals supplied only paediatric, or only adult recordings. Therefore, we implemented a linear mixed model (LMM) in each frequency band, specifically a random intercept model, with the originating hospital site as a random effect. Our cohort has a wide range of subjects per hospital (Figure 1), but one of the strengths of mixed modelling is the ability to handle unequal group sizes (West et al., 2014).
Possible fixed effects were age and sex, meaning four LMMs were under consideration in each frequency band:
Null: $RBP(\cdot)\sim(1|Hospital)$
Age: $RBP(\cdot)\sim Age+(1|Hospital)$
Sex: $RBP(\cdot)\sim Sex+(1|Hospital)$
Full: $RBP(\cdot)\sim Age+Sex+(1|Hospital)$
An interaction term between age and sex was also considered but deemed unnecessary (Supplementary S7). Model fitting could be performed at either the regional or whole-brain level. Taking the latter, we calculated several model evaluation statistics to determine the optimal fixed effect structure of the LMM in each frequency band. Since we worked at the whole-brain level, the original data table was used. The statistics under consideration were the AIC and BIC for each model, 95% profiled confidence intervals for the regression coefficients of fixed effects in all but the null model (which has none) and the likelihood ratio test $p$-values for every pair of models which differed by one variable.
In four frequency bands, each metric was in agreement regarding the optimal LMM, so no further tests were required. The preferred models were the age model in the $\delta$ and $\beta$ bands; the sex model in $\gamma$ and the full model in $\theta$. The $\alpha$ band demonstrated some uncertainty between the age model and the full model. Since both models perform similarly, we retain the simpler age model as the optimal choice in the $\alpha$ band.
Hence, the optimal LMM was frequency band-specific in our cohort, and so we report frequency band-specific effects throughout our results. We have provided a brief visual overview of our methods (Figure 2). In the results section, we explore both random (hospital site) and fixed (age, sex) effects. Specifically, we investigate if the variables explain RBP($\cdot$) variation, and examine how they might influence it.
Figure 2: A visual representation of our methods: electrode localisation (top) and signal processing following segment selection (middle), for each subject. Subjects were combined to create our normative data table, to which we applied our statistical models (bottom). All components show dummy data for example purposes only. Brain plots from Scholtens et al. (2021).
@@ 3 Results
3.1 Hospital site effects impacted icEEG normative maps
Focusing first on the random effect structure (recording hospital), we found that in our cohort and model, the hospital site effect was much more powerful in explaining RBP($\cdot$) differences at the whole-brain level than any fixed effect structure, in almost all cases. Table 1 quantifies this, showing marginal $R^{2}$ values ($R^{2}_{m}$) and intraclass correlation coefficients ($ICC$) as percentages for the full model, the age model and the sex model. The $R^{2}_{m}$ represents the proportion of variation in RBP($\cdot$) explained by the fixed effect(s) of that model. The $ICC$ represents the proportion of the variance explained by the recording hospitals. The optimal LMM as per Section 2.5 has been highlighted for each frequency band.
Note, the $R_{m}^{2}$ values of the two single fixed effect models do not sum exactly to the $R^{2}_{m}$ of the full model, because the denominator in the calculation of $R_{m}^{2}$ includes the residual variance and therefore changes with each model.
@@ 3.1 Hospital site effects impacted icEEG normative maps
$\delta$ 4.28 15.23 4.27 15.27 0.00 19.42
$\theta$ 0.60 15.17 0.23 14.88 0.41 16.25
$\alpha$ 8.04 5.09 7.97 5.13 0.00 5.88
$\beta$ 2.59 6.58 2.58 6.58 0.00 7.93
$\gamma$ 0.25 32.59 0.01 32.16 0.24 32.25
Table 1: $R^{2}_{m}$ and $ICC$ values (measured in %) for the full model, the age model and the sex model. $R^{2}_{m}$ represents the proportion of variation in RBP($\cdot$) explained by the fixed effect(s) of that model. The $ICC$ represents the proportion of the variance explained by the grouping structure, namely recording hospitals. The optimal covariate subset, as determined by a standard model selection process, is highlighted for each frequency band.
There is striking variation in the magnitude of the hospital site effect on RBP($\cdot$) across signal properties. However the $ICC$s consistently exceed 5%, indicating that, in our cohort, at least a twentieth of the variation in RBP($\cdot$) was explained by the random effect structure. This proportion sometimes reached as high as 30%, attributing that recording hospital impacted RBP($\cdot$) in all frequency bands. Further, in all bands except $\alpha$, the $ICC$ was consistently larger than the $R^{2}_{m}$ values across models, with this difference being most notable in the $\theta$ and $\gamma$ bands.
To visualise the effect of the originating hospital on RBP($\cdot$), we fit the age model to all data, then selected three well-populated hospitals with similar age ranges and plotted RBP($\delta$) against age, along with the model fit in Figure 3. Additionally, we identified a 33-year-old male from each of the three hospitals and have provided the first 10 seconds of their processed 70-second icEEG segment. The three subjects had variable numbers of channels, so only the first 50 channels are displayed for comparative purposes. Figure 3 demonstrates that while individuals’ icEEG segments may appear similar, the originating hospital introduces underlying differences to relative band power properties. Visualisations for all five frequency bands and all hospitals can be found in Supplementary S8.
Figure 3: Scatterplot of RBP($\delta$) against age for thee well-populated hospitals. The age model was implemented for all hospitals, then the age slope and intercepts for Jefferson Hospital, Mayo Clinic and University College London Hospital were extracted and plotted. For three similar individuals, men of age 33, the first 50 channels and 10 seconds of their 70-second preprocessed icEEG segments have been provided. The grey bar indicates where the segments from the 33-year-old men lie within this subset of the data.
@@ 3.2 Age and sex explained only some of the variation in relative band power
We next examined our fixed effects at the whole-brain level ($R^{2}_{m}$). Neither age nor sex explained a substantial portion of the variance in the RBP($\cdot$), in any frequency band (Table 1). The largest effect was found in the $\alpha$-band, where $R^{2}_{m}$ was highest at 8%, which in contrast, is similar to the minimum $ICC$ values at around 5%.
@@ 3.3 Age was more important than sex for predicting relative band power
Comparing $R^{2}_{m}$ for the age and sex models revealed which of the two was more valuable when predicting RBP($\cdot$). All bands which achieve $R^{2}_{m}>1\%$ do so only when age is included in the model (Table 1). In fact, sex consistently accounted for a negligible portion of the variance in the response, with $R^{2}_{m}$ values below 0.5%. This statement holds even in $\theta$ and $\gamma$, with sex retained in the optimal LMM, confirming sex was not a significant predictor in this cohort. In contrast, $R^{2}_{m}$ values for age fluctuated notably across frequency bands, being higher in $\delta$, $\alpha$ and $\beta$ ($R^{2}_{m}$ range of $7.96\%$ and $0.41\%$ for age and sex models respectively).
To validate the lack of sex effect in this cohort, we plotted RBP($\cdot$) against age for all frequency bands at the whole brain level. We then fit the full model and plotted the male/female regression lines. Figure 4 demonstrates visually that sex is not a significant predictor in this cohort, with the differing sex lines being indistinguishable in almost all bands.
Together, Table 1 and Figure 4 emphasize that the influence of sex was minor, and age was the more important predictor of RBP($\cdot$) in our cohort. Hence, subsequent analysis focuses on the age model for all frequency bands; however, we exercise caution when interpreting results in the $\theta$ and $\gamma$ bands, which have a weak relationship with both fixed effects.
Figure 4: Scatterplots of RBP($\cdot$) against age in every frequency band, using data at the whole-brain level. The full model has been implemented and the male/female regression lines have been plotted.